# Multimodal Fusion on Low-quality Data: A Comprehensive Survey

### v2

@@ 5.4 Discussion
Dynamic multimodal learning methods focuse on the variance of modality quality with respect to sample, time or space, which widely exists but tends to be neglected. The dynamic fusion methods include heuristic (mostly designed for specific applications), attention-based (typically for representation fusion) and uncertainty-aware (modeling the modality and sample level uncertainty for fusion) strategies. There are great potentials for dynamic multimodal learning. First, the dynamic principle could be considered in SOTA multimodal models, e.g., CLIP. Second, there are lots of dynamic scenarios in real applications (e.g., autonomous driving, medical image fusion), so designing application-specific dynamic fusion strategies is interesting. For example, in multimodal medical images, we can dynamically fuse them at path-level, which may providing better flexibility and interpretablility.

### v3

@@ 5.4 Discussion
Dynamic multimodal learning methods focuse on the variance of modality quality with respect to sample, time or space, which widely exists but tends to be neglected. The dynamic fusion methods include heuristic (mostly designed for specific applications), attention-based (typically for representation fusion) and uncertainty-aware (modeling the modality and sample level uncertainty for fusion) strategies. There are great potentials for dynamic multimodal learning. First, the dynamic principle could be considered in SOTA multimodal models, e.g., CLIP. Second, there are lots of dynamic scenarios in real applications (e.g., autonomous driving, medical image fusion), so designing application-specific dynamic fusion strategies is interesting. For example, in multimodal medical images, we can dynamically fuse them at path-level, which may providing better flexibility and interpretability.